User Tools

Site Tools


handover:hob008

HOB008

Hobart 26m Disk VSN:HOB+0126

Data Volume at beginning: 7235 GB

  • 2000UT Data not being recorded initially, reloaded schedule and procedure files (Ross)
  • 0055 - Restarted DBBC Control program as clkoff/dbbc=pps_delay readings indicated a loss of synch. (JMc).
  • Missed scans 324-0420, 324-0417 and 324-0422 due to 'previous source in the schedule not reached before new source was commanded …'. Alarm did not go off for this (Lucas)
  • 0830UT - Disc pos ~200 GB behind, holding steady. (Ellen)
  • 0845UT - Hg pressure still at about 4.0 (Lucas reported it had been a little high). Appears stable, probably just due to hot weather. Continuing to monitor. (Ellen)
  • 0900UT - Hg pressure fallen down to 3.1 microns. (Ellen)
  • 1040UT - Module filled up and automatically switched to recording onto HOB+0120, which already had 9901 GB on there. First scan on module was 324-1041. 2092 GB recorded to HOB+0029. (Ellen)

Hobart 12m

Disk VSN: HOB+0029

Data Volume at beginning: 3577 GB

  • 1900UT Experiment started OK (Ross)
  • 22:30 to 23:06UT, getting mk5cn time-out connection closed errors, restarted dimino and dbbc (and re-synced) but continued getting errors. Called on-call person, they recommended editing checkmk5 file in pfmed to exclude /sys/… command which produced autocorrelations by commenting out with “, but this didn't work either. Re-started schedule as the error was deemed to be benign. Alarm was commented out of alarm_control (Lucas/Ross)
  • 0215 UT The diskpack that was being recorded to was marked as 'Faulty', and there was a lag between accepting a 'disk_record=off' command, which was longer than the 500ms window the fs allowed, thereby giving the mk5cn time-out error. Jamie was able to correct this and the alarm was un-commented. (Lucas/Jamie)
  • 0830UT - Disc pos ~130 GB behind, holding steady. (Ellen)
  • 1040UT - Module filled up and automatically switched to recording onto HOB+0120, which already had 9901 GB on there. First scan on module was 324-1041. 2092 GB recorded to HOB+0029. (Ellen)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/hob008.txt · Last modified: 2016/11/19 12:40 by Ross Turner